If x+a=(b/3)x and b=/ 3, then x =?

Answer by stanbon(75887)

If x+a=(b/3)x and b is not = 3, then x =?
---
x - (b/3)x = -a
x(1-(b/3)) = -a
---
x = -a/[1-(b/3)]
---
x = -a/[(3-b)/3]
---
x = -3a/(3-b)
=================
Cheers,
Stan H.